OpenStudy (usukidoll):

Sometimes, but not always, quantifiers distribute over logical operations. Determine which of the following pairs of statements are equivalent. In the case of nonequivalent pairs, give an example of propositional functions P(x) and Q(x) for which the paired statements are not equivalent problems and attempts coming in a bit...

OpenStudy (usukidoll):

a. (∀x) [P(x) ^ Q(x)] and (∃x)P(x) ^ (∃x)Q(x) I got. for the left hand side. All x's with p and q. For the right hand side Some x's with p or some x's with q. Is there a way to tell that they're not equivalent or are equivalent ? What if I switch P(x) and Q(x) in these problems? Or let P(x) and Q(x) be a function

OpenStudy (usukidoll):

b. (∃x)[P(x) ^ Q(x)] and (∃x)(P(x)) ^ (∃x)(Q(x)) left: some x's with p and q right: some x's with p and some x's with q

OpenStudy (usukidoll):

c.(∀x)[P(x) V Q(x)] and (∀x)P(x) V (∀x)Q(x) left There exists all values of x with p or q right There exists all values of P or there exists all values of Q.

OpenStudy (usukidoll):

d.(∃x)[(P(x) V Q(x)] and (∃x)P(x) V (∃x)Q(x) There exist some values of x in P or Q There exist some values of P or there exists some values of Q

OpenStudy (usukidoll):

( ∀x) [P(x) ---> Q(x) ] and ( ∀x)P(x) ---> ( ∀x)Q(x) Left: There exists all values of x in P as well as Q. Right: All x values in P will also have all x values in Q

OpenStudy (usukidoll):

f. ( ∀x)[P(x)<->Q(x)] and ( ∀x)P(x) <-> ( ∀x)Q(x) All x values are in p if and only if there's q All x values are in p if and only if there are all x values in q.

OpenStudy (usukidoll):

@wio @UnkleRhaukus @inkyvoyd

OpenStudy (usukidoll):

∀ - for all ∃ - for some ∃! - there is one unique solution

OpenStudy (usukidoll):

@thomaster

OpenStudy (anonymous):

are you asking whether the two are equivalent?

OpenStudy (usukidoll):

yeah

OpenStudy (usukidoll):

first you have to write it out to get a better understanding... then you have to determine if they are equivalent...

OpenStudy (anonymous):

I don't think the two are equivalent.

OpenStudy (usukidoll):

ok how can we prove that?

OpenStudy (usukidoll):

but moreover, is the sentence structure correct to begin with?

OpenStudy (unklerhaukus):

I dont think they are equivalent either, the easiest method to disprove , would be to find a single counter example

OpenStudy (anonymous):

i was thinking if you can find an x the P(x) and Q(x) is not true, then the statement on the right side is true, but the one on the left side is false

OpenStudy (usukidoll):

like we could have P(x) be something. Q(x) be something as well.

OpenStudy (anonymous):

this is a tough one....

OpenStudy (usukidoll):

and then apply to the...........problems... maybe :/

OpenStudy (usukidoll):

(∀x) [P(x) ^ Q(x)] and (∃x)P(x) ^ (∃x)Q(x) means For all of the x's P and Q. but the right one is For some x, there is P and for some x there is q

OpenStudy (usukidoll):

maybe we could let P(x) and Q(x) be a function... ?

OpenStudy (anonymous):

we could. P(x) and Q(x) can represent anything

OpenStudy (usukidoll):

like let P(x) = 3x+ y = 8 hahah I'm making this up Q(x) = 4x+ 6 = 10 ???

OpenStudy (anonymous):

what about y?

OpenStudy (usukidoll):

oh sdfjkaljfdl; so make Q(x) = 4x + 6y =10

OpenStudy (anonymous):

The problem I was thinking is that does x have to be the same object in the statement (∃x)P(x) ^ (∃x)Q(x)

OpenStudy (usukidoll):

I know that it's for some x there is P and for some x there is Q

OpenStudy (anonymous):

(∃x)P(x) ^ (∃ *anotherP x)Q(x)?

OpenStudy (anonymous):

(∃x)P(x) ^ (∃ *another x)Q(x)?

OpenStudy (anonymous):

typo :D

OpenStudy (usukidoll):

like a y?

OpenStudy (usukidoll):

@agent0smith torture time!

OpenStudy (anonymous):

yea, that shows that x and y are not necessarily a same object

OpenStudy (usukidoll):

hmmmmm I don't get it :/ unless (∃x)P(x) ^ (∃ y)Q(x)?

OpenStudy (usukidoll):

so for some x there is a p and for some y there is a q

OpenStudy (usukidoll):

that doesn't say much unless you make P(x) and Q(x) a simple function or something . or sdfljkad;k

OpenStudy (anonymous):

(∃x)P(x) ^ (∃ y)Q(y)

OpenStudy (usukidoll):

for some value of x there is p and for some value of y there is q

OpenStudy (usukidoll):

except that we have x and y coming into play

OpenStudy (usukidoll):

@Isaiah.Feynman

OpenStudy (isaiah.feynman):

Not here yet. What maths is this?

OpenStudy (anonymous):

let x be a member in [-2,2], let p(x) := sqrt(1-x^2) ≥ 0 , and Q(x) = x^2 ≥ 0 then, (∃x)P(x) ^ (∃x)Q(x) is true, but, (∀x) [P(x) ^ Q(x)] is false

OpenStudy (usukidoll):

alright so .. for some x there is p and for some x there is q. so we can apply a number into those functions.

OpenStudy (anonymous):

yep, that is a good counter example :DD

OpenStudy (usukidoll):

like let x = 2 p(2) = sqrt ( 1-4) ≥ 0 Q(2) = 4 ≥ 0

OpenStudy (usukidoll):

:O that's a very big number . I don't see how sqrt(-3) can be greater than 0.

OpenStudy (usukidoll):

maybe -2 would work better.

OpenStudy (usukidoll):

P(-2) sqrt(1-(-2)^2) >0 Q(-2) = 4 ≥ 0

OpenStudy (agent0smith):

Not read the whole question but the only way p(x) := sqrt(1-x^2) ≥ 0 is true is if x =< 1

OpenStudy (usukidoll):

oh there is some value... some values of x may work to satisfy p and q. but not all of the values of x would work.

OpenStudy (usukidoll):

(∀x) [P(x) ^ Q(x)] is all values of x in P and Q but if let P and Q be simple equations and test out all the values of x, something is going to be false. either P(x) or Q(x) maybe both will give out false results.

OpenStudy (usukidoll):

how can 0 be greater than or equal to 0 ?eh that' s 0 = 0

OpenStudy (agent0smith):

0 is greater than or EQUAL to zero :P

OpenStudy (usukidoll):

let x = fractions :D

OpenStudy (anonymous):

again, we're trying to disprove (∀x) [P(x) ^ Q(x)] and (∃x)P(x) ^ (∃x)Q(x) are equivalent by find a counter example

OpenStudy (anonymous):

which i did

OpenStudy (usukidoll):

that wasn't me who did that 0 thing.. but you did mention that let x [-2,2]

OpenStudy (usukidoll):

"let x be a member in [-2,2], let p(x) := sqrt(1-x^2) ≥ 0 , and Q(x) = x^2 ≥ 0 then, (∃x)P(x) ^ (∃x)Q(x) is true, but, (∀x) [P(x) ^ Q(x)] is false"

OpenStudy (usukidoll):

so I pick any number between -2 and 2 that produces non equivalency or equivalency

OpenStudy (agent0smith):

p(x) := sqrt(1-x^2) ≥ 0 is true is if -1 =< x =< 1 and Q(x) = x^2 ≥ 0 true if x >= 0 so 0 =< x =< 1

OpenStudy (usukidoll):

0 is a BAD NUMBER D:

OpenStudy (usukidoll):

so for some values of x, the conditions of p and q are satisfied... but not values of x can do that.

OpenStudy (usukidoll):

and I want a pizza now... T_T

OpenStudy (anonymous):

well, to show that (∃x)P(x) ^ (∃x)Q(x) is true, let x = 0 sqrt(1-0^2) ≥ 0 is true, and 0^2 ≥ i strue. however, P(x) and Q(x) are not true for all x in [-2,2]. let x = -2

OpenStudy (anonymous):

typo 0^2 ≥0 is true

OpenStudy (usukidoll):

:S 0^2 is greater than or equal to 0... how D:

OpenStudy (anonymous):

0 ≥ 0 is a true statement

OpenStudy (anonymous):

remmeber, ≥ means greater OR equal

OpenStudy (agent0smith):

0 is equal to 0. Therefore it is greater than OR equal to zero.

OpenStudy (usukidoll):

oh rightttttttttt...

OpenStudy (usukidoll):

so do I come up with my own P(x) and my own Q(x) for each of these situations?

OpenStudy (usukidoll):

@satellite73 I know you're here ^^

OpenStudy (anonymous):

yeah, you can come up with a different counter example. But the important thing is that do you now see that (∀x) [P(x) ^ Q(x)] and (∃x)P(x) ^ (∃x)Q(x) are not equivalent?

OpenStudy (usukidoll):

yes I could... if I come up with my own p(x) and q(x) to prove that they're not equivalent by letting x be a number... it's easier to see how these conditions can easily be broken.

OpenStudy (usukidoll):

so I create my own P(x) function and q(x) function for (∃x)[P(x) ^ Q(x)] and (∃x)(P(x)) ^ (∃x)(Q(x)) That states that for some x there is p and q and on the right there is some x for p and some x for q

OpenStudy (usukidoll):

so let P(x) = 5x + 2 = 7 lol maybe >:) Q(x) = 4x + 1 = 5 XD x = 1 that might work for for some x there is p and q

OpenStudy (anonymous):

well, you need to state what your domain is

OpenStudy (usukidoll):

domain?

OpenStudy (usukidoll):

umm [-1,1] ?

OpenStudy (anonymous):

that works

OpenStudy (usukidoll):

so for this we do have some value of x that works for p and q and some value of x that works for p and some value of x that works for q while x = -1. 0 doesn't work, x = 1 works :)

OpenStudy (anonymous):

yes, 1 works. Everything else in [-1,1] doesn't

OpenStudy (usukidoll):

c.(∀x)[P(x) V Q(x)] and (∀x)P(x) V (∀x)Q(x) left There exists all values of x with p or q right There exists all values of P or there exists all values of Q. So P(x) and Q(x) must be a function that satisfies all values of x. ANd since this is an or statement, it shouldn't be that hard to test for equivalency

OpenStudy (usukidoll):

:S I found a typo in my A T___T gawd I feel horrible ... the left side is correct but the right is there exists all values of x for p or there exists all values of x for q. D*****!*!*!(&#(*!&@(*!&(*

OpenStudy (usukidoll):

(∀x) [P(x) ^ Q(x)] and (∀x)P(x) ^ (∀x)Q(x) changes everything!

OpenStudy (usukidoll):

d.(∃x)[(P(x) V Q(x)] and (∃x)P(x) V (∃x)Q(x) There exist some values of x in P or Q There exist some values of P or there exists some values of Q Let the domain be [-1.1]. Moreover let P(x) = x+ 2 = 3 and Q(x) = x-3 = -2 If x = 1 then it satisfies P(x) and Q(x).

OpenStudy (usukidoll):

oops it satisfies P(x) or Q(x) for d. oy

OpenStudy (usukidoll):

(∀x) [P(x) ^ Q(x)] and (∀x)P(x) ^ (∀x)Q(x) is A. UGH! That means .. for all values of x in P and Q AND for all values of x in P and for all values of x in Q. so all the x's out there need to satisfy or we can do a contradiction.

OpenStudy (usukidoll):

hmmm the domain needs to be a bit bigger for d. because this time we can satisfy p OR Q

OpenStudy (usukidoll):

unless we make P(x) x^2 >0 Q(x) x < 0 from [-1.1] yES! That satisfies one or the other.

OpenStudy (usukidoll):

d.(∃x)[(P(x) V Q(x)] and (∃x)P(x) V (∃x)Q(x) There exist some values of x in P or Q There exist some values of P or there exists some values of Q Let the domain be [-1.1]. Moreover, let P(x) = x^2 >0 and Q(x) = x <0 If we let x = -1, it doesn't satisfy P, but it satisfies Q If we let x = 1, it satisfies P, but it doesn't satisfy Q. SInce we have an or statement, we can choose to satisfy P or Q.

OpenStudy (usukidoll):

oh wait try... P = x > 0 and Q = x < 0 domain from -1,1

OpenStudy (usukidoll):

@sourwing

OpenStudy (usukidoll):

or statements might be easier to do... I just thought of letting P(x) = x^2 > 0 and Q(x) = x^3 > 0 :O I think I figured out my A problem. for all values of x so let the domain be [-2,2] and plug into these equations P(x) = x^2 > 0 and Q(x) = x^3 + 9 > 0 it works! because all of the results are greater than 0

OpenStudy (usukidoll):

(∀x) [P(x) ^ Q(x)] and (∀x)P(x) ^ (∀x)Q(x) for all values of x there is p and q and for all values of x there is p and for all values of x there is q. THat example alone satisfies this ^

OpenStudy (usukidoll):

@LastDayWork wassup ^^